OpenStudy (anonymous):

How do i find the number of terms in geometric sequence: 1/2, 1/4, 1/8, ..., 1/1024?

OpenStudy (anonymous):

cos I got 19.54 = 1/2^n-1 and I know it's wrong D:

OpenStudy (anonymous):

tn=ar^(n-1) 1/1024 = .5(.5^n-1) 1/512=.5^n-1 log 1/512 = log .5^n-1 -2.71= (n-1) log .5 -2.71=(n-1) (-.3) 9=n-1 10=n

OpenStudy (anonymous):

oohh, thank you!! :D

OpenStudy (anonymous):

nytime